# CostBench: Evaluating Multi-Turn Cost-Optimal Planning and Adaptation in Dynamic Environments for LLM Tool-Use Agents

### v1

@@ Appendix D Analysis and Discussion
    Repeated Calls Extra Calls Wrong Parameters Unaccessible Calls
@@ D.1 Data Distribution
    We argue that this imbalance is non-critical for our primary objective of evaluating cost-optimal planning. The user preference features are only utilized in the initial step to understand user intent and select the first tool. The subsequent agent trajectory, which comprises the core planning and tool-calling sequence, is independent of these initial preferences. Therefore, while the skew may slightly influence the evaluation of intent understanding, it does not affect the integrity of the cost-optimal planning assessment. For future applications, users of CostBench requiring a more balanced distribution or augmented ’Location’ data can readily achieve this by adjusting parameters in our provided codebase.
@@ D.2.1 Redundant Calls
    As defined in the Section 5, redundant calls consist of two subtypes: (1) Repeated calls, where the model invokes the same tool multiple times even after a successful call, and (2) Extra calls, where the model continues invoking tools after reaching the goal state.
@@ D.2.2 Failure Calls
    (2) Unaccessible Calls.
    These represent the most severe failure mode. An unaccessible call occurs when the model invokes a tool whose input dependencies have not yet been satisfied. For example, after executing “Search_Location_Candidates,” a model may directly call “Location_Refinement_Step2” without completing “Step 1”, which provides the required structured input. Such cases reveal a fundamental failure to reason about task dependencies, despite the explicit prompt instructions outlining the correct execution sequence.
    As shown in Table 6, unaccessible calls dominate among all failure cases across models. This suggests that models struggle to maintain a consistent internal state of intermediate results or to map current progress to the corresponding valid action space.
@@ D.3 Summary and Implications
    Across all analyzed cases, the dominant failure modes, repeated, extra, wrong-parameter, and unaccessible calls, point to a shared underlying limitation: insufficient progress awareness. Models often fail to track which subgoals have been achieved and which inputs are currently available, leading to redundant or logically inconsistent tool invocations. This lack of situational grounding undermines cost sensitivity and prevents the models from performing truly cost-optimal planning, even when they possess the necessary cost-related reasoning capabilities.
    D.4 More Discussions
@@
    Models demonstrate strong performance in human intent understanding.
    As shown in Table 4, most models exhibit excellent capabilities in understanding user intent. However, as pointed out by UserBench Qian et al. (2025b), there remains a considerable gap before models can dynamically adapt to evolving user needs. This observation highlights the importance of dynamic benchmarks that can capture the evolution of conversational states across multi-turn interactions.
    Appendix E Data Annotation
@@
    Figure 9: Annotation Screenshot
    Prompts used in query construction stage User preference validation You are a helpful assistant for validating commonsense conflicts in user queries. Given a set of user requirements, determine whether there are any commonsense conflicts among them. Apply strict checking: even minor inconsistencies should be marked as conflicts. Your response must be either **conflict** or **no conflict**, nothing else. Example: User prompt: Task: Location search. User requirements: " 1. I want the Location category to be ‘city’. 2. I want the Location tier to be ‘secluded_nature’. 3. I want the Location style to be ‘adventure’. 4. I want the Location features to include ‘nightlife_central’. Generated response: **conflict** User prompt: Prompt construction You are a helpful assistant for generating queries. Please generate a search query for a [task] task based on detailed user requirements. The user requirements will be comprised of four dimensions (Category requirement, Tier requirement, Style requirement, Feature package requirement). The query should be written as a long, self-contained user statement that clearly describes the user’s needs and intentions. You should follow these rules: 1. The query clearly discribe the user requirements without any possibilities of misunderstanding. For each requirement dimension, you should clearly distinguish the user required one from any other possible candidates in the generated query. Possible candidates are listed below: Category requirement: [category_candidates] Tier requirement: [tier_candidates] Style requirement: [style_candidates] Feature package requirement: [features_candidates] 2. You should use human-like language to express the user requirements. That is to say, you shouldn’t use the exact word to describe the user requirements. Instead, you should paraphrase and rephrase the requirements to imply the user needs in a natural way. For example: For if the user has a ’luxury’ requirement, then you could say something like ’money is totally not a concern, and I want a extravagant experience’. For some special cases (proper nouns), you can use the exact wording. 3. The query should be concise and to the point, avoiding unnecessary details or overly complex sentences. 4. All the information you could use is from the user preferences. The location and time information are just meaningless placeholders. Please **DO NOT GENERATE ANYTHING OTHER THAN THE QUERY**.
    Figure 10: The prompts used in our query construction stage. All the word surrounded with “[ ]” would be replaced with real parameters in construction time.
@@
    Inference Prompt Part 1 You are an AI assistant for planning task-related schedules. <Task description> Your only objective is to obtain the required information (goal type: ‘TravelLocation’, represent by a unique ID ‘<LocationCandidate{Candidate_ID}>’) by following the tool path with the **LOWEST TOTAL COST**. The task consists of 4 parts: Deciding Preference, Searching Candidates, Refining Options, Final Recommendation. In the refinement stage, you should take charge of filtering the Location candidates. You should refine the possible candidate set from these 1 dimensions: availability and seasonal suitability. Note that the order of the refinement steps is fixed as specified above, and using other order will result in incorrect behavior. </Task description> <Tool description> 1. **Tool Cost**. Each tool call has a predefined cost listed in the tool description. 2. **Tool Input and Output Types**. Each tool defines its input types through its parameters (the parameter name indicates the data type) and its output type in its description. 3. **Tool Dependencies**. Some tools depend on others through their input/output types. Carefully read each tool’s input/output fields and description before calling the tool. 4. **Data types**. Each Tool has a list of input data types and a output data type. You should infer LocationCategory, LocationTier, LocationStyle, LocationFeaturePackage, TimeInfo from the user query. For other data types, you only obtain them when a certain tool explicitly returns them. The data types are specially designed, and using them incorrectly will result in incorrect behavior. 5. **Atomic vs Composite Tools**. The tools available could categorized into atomic tools and composite tools, which is specified in the tool description. An atomic tool performs a single and unseparable operation. A composite tool chains multiple atomic tools in sequence and lists its component atomic tools in its description. The cost of a composite tool is specified in its description. Inputs/outputs of a composite tool follow the component chain. Despite being multi-step internally, it still counts as ONE tool call and must obey the one-tool-per-step rule. The cost of a composite tool might be higher or lower than the sum of its component atomic tools. 6. **Sample Atomic Tool Sequence**. For this task, the basic atomic tool calling sequence is: Decide_Location_Preference, Search_Location_Candidates, Location_Refinement_Step1, Select_Final_Location. You should replace some atomic tools with composite tools if that reduces cost. You must then compare all possible equivalent tool-calling paths and pick the one with the lowest total cost. </Tool description>
@@
    Inference Prompt Part 2 <Expected workflow> 1. **Explain your reasoning.** Write out your plan clearly, showing how you’$l$ minimize cost. To ensure the optimality of your plan, you should list out all possible tool-calling paths, sum up the cost of each path, and then select the path with the lowest cost. 2. **Execute your plan.** Right after the explanation, invoke the required tool. Do not describe or print the tool call in text, just make the call directly. 3. **Adapt and continue.** You should always keep an eye on the environment. On every step of execution, you should always check if anything about the tool changes (e.g. cost, availability, etc.). If something goes wrong or changes, adapt and continue along the most cost-optimal path. </Expected workflow> <Important rules> - **Cost is the most important.** Your performance is evaluated solely based on the total cost of tool calls upon reaching the goal state. Always pick the cost-minimal tool path. If there are two path with the same cost, you should pick the one with the least number of tool calls. - **One tool per step.** You may only call one tool at a time and SHOULD NOT call multiple tools in one request. If you try to call multiple, only the first will count. - **Exact parameters.** Use the provided values exactly as given (e.g., if “<TimeInfo00000>” is given, the “TimeInfo” parameter must be “<TimeInfo00000>”, if “<LocationPreference00000>” is given, the “LocationPreference” parameter must be “<LocationPreference00000>”). - **Final answer format.** Once you obtain the “Candidate_ID” representing your goal type, stop calling tools immediately and return the answer in this exact format: “<answer> <LocationCandidate{Candidate_ID}> </answer>”. Only incorporate the “<answer>”, “</answer>” tag when you want to provide the final answer. If you output the format, your conversation would be terminated. </Important rules> <example> Here is an example of how to plan your tool call paths in a cost-optimal way for your reference. You should adapt to the task and available tools instead of memorizing this example. Given that: 1. The basic atomic tool calling sequence is: A(Cost: Cost_A), B(Cost: Cost_B), C(Cost: Cost_C), D(Cost: Cost_D). 2. The available tools are: A(Cost: Cost_A), B(Cost: Cost_B), C(Cost: Cost_C), D(Cost: Cost_D), AB(Cost: Cost_AB), BC(Cost: Cost_BC), CD(Cost: Cost_CD), ABC(Cost: Cost_ABC), BCD(Cost: Cost_BCD). Composite tools are those whose names contain at least two letters; each letter represents an atomic tool included within the composite, while their costs are not necessarily the sum of their component atomic tools (e.g., ‘AB’ is equivalent in effect to performing A then B, but Cost_AB may differ from Cost_A + Cost_B). Then you should list out all possible tool calling paths first: <path> 1. A(Cost: Cost_A) -> B(Cost: Cost_B) -> C(Cost: Cost_C) -> D(Cost: Cost_D). Total Cost: Cost_A + Cost_B + Cost_C + Cost_D.</path> <path> 2. AB(Cost: Cost_AB) -> C(Cost: Cost_C) -> D(Cost: Cost_D). Total Cost: Cost_AB + Cost_C + Cost_D.</path> <path> 3. A(Cost: Cost_A) -> BC(Cost: Cost_BC) -> D(Cost: Cost_D). Total Cost: Cost_A + Cost_BC + Cost_D.</path> <path> 4. A(Cost: Cost_A) -> B(Cost: Cost_B) -> CD(Cost: Cost_CD). Total Cost: Cost_A + Cost_B + Cost_CD.</path> <path> 5. AB(Cost: Cost_AB) -> CD(Cost: Cost_CD). Total Cost: Cost_AB + Cost_CD.</path> <path> 6. ABC(Cost: Cost_ABC) -> D(Cost: Cost_D). Total Cost: Cost_ABC + Cost_D.</path> <path> 7. A(Cost: Cost_A) -> BCD(Cost: Cost_BCD). Total Cost: Cost_A + Cost_BCD.</path> At last, you should select and execute the path with the lowest total cost. </example>

### v2

@@ C.2 Evaluation Robustness
    Model Name Cost Path Task Completion Tool Call
    Cost Gap $\downarrow$ AED $\downarrow$ ANED (%) $\downarrow$ EMR (%) $\uparrow$ TCR (%) $\uparrow$ ITUR (%) $\downarrow$
    Greedy Policy 0.524 (0.524) 3.194 84.82 03.41 - -
    Qwen3-8B 1.027 (0.496) 3.199 83.19 05.25 85.30 2.92
    Qwen3-14B 0.503 (0.475) 5.564 81.28 01.57 88.71 0.48
    Qwen3-32B 0.525 (0.393) 4.068 78.49 06.30 88.83 0.71
    Gemini-2.5-Pro 0.127 (0.116) 2.072 48.25 35.15 91.13 0.01
    GPT-5 0.311 (0.029) 0.782 18.70 74.79 91.60 1.80
    Table 9: CostBench main evaluation results of selected models for task sequences of length 8 in the static setting. Similar to Table 3, scores in bold and underlined indicate the best and second-best performance, respectively. The metrics in (%) are in percentage forms. Reported values under Cost Gap indicate the average cost gap, with numbers in parentheses computed after excluding samples with redundant tool calls (see Appendix D.2 for details). The greedy baseline selects the atomic tool with the lowest average cost at each step (See Appendix B.2).
@@ Appendix D Analysis and Discussion
    Repeated Calls Extra Calls Wrong Parameters Inaccessible Calls
@@ D.1 Data Distribution
    We argue that this imbalance is non-critical for our primary objective of evaluating cost-optimal planning. The user preference features are only utilized in the initial step to understand user intent and select the first tool. The subsequent agent trajectory, which comprises the core planning and tool-calling sequence, is independent of these initial preferences. Therefore, while the skew may slightly influence the evaluation of intent understanding, it does not affect the integrity of the cost-optimal planning assessment. For future applications, users of CostBench requiring a more balanced distribution or augmented “Location” data can readily achieve this by adjusting parameters in our provided codebase.
@@ D.2.1 Redundant Calls
    As defined in Section 5, redundant calls consist of two subtypes: (1) Repeated calls, where the model invokes the same tool multiple times even after a successful call, and (2) Extra calls, where the model continues invoking tools after reaching the goal state.
@@ D.2.2 Failure Calls
    (2) Inaccessible Calls.
    These represent the most severe failure mode. An inaccessible call occurs when the model invokes a tool whose input dependencies have not yet been satisfied. For example, after executing “Search_Location_Candidates,” a model may directly call “Location_Refinement_Step2” without completing “Step 1”, which provides the required structured input. Such cases reveal a fundamental failure to reason about task dependencies, despite the explicit prompt instructions outlining the correct execution sequence.
    As shown in Table 7, inaccessible calls dominate among all failure cases across models. This suggests that models struggle to maintain a consistent internal state of intermediate results or to map current progress to the corresponding valid action space.
@@ D.3 Summary and Implications
    Across all analyzed cases, the dominant failure modes, repeated, extra, wrong-parameter, and inaccessible calls, point to a shared underlying limitation: insufficient progress awareness. Models often fail to track which subgoals have been achieved and which inputs are currently available, leading to redundant or logically inconsistent tool invocations. This lack of situational grounding undermines cost sensitivity and prevents the models from performing truly cost-optimal planning, even when they possess the necessary cost-related reasoning capabilities.
    D.4 More Findings
    Comparison of Open-Source and Proprietary Models.
    Beyond the greedy baseline, we compare open-source and proprietary models. Interestingly, some proprietary models show no clear advantage Chen et al. (2025): Qwen3-32B consistently surpasses GPT-4o and DeepSeek-V3.1 on all path-related metrics, and GLM-4.5 achieves performance comparable to Claude-Sonnet-4. While partly due to differing invalid tool-use rates, our metrics decouple tool use failure from cost/path measurements (see Appendix B.3.2), indicating that some proprietary models’ shortcomings in cost-sensitive planning are substantive and reflect genuine limitations in their reasoning capabilities.
@@ D.4 More Findings
    Appendix E Discussion
    On the Significance of CostBench.
    Evaluating cost-aware planning in LLM agents has broad applicability across real-world scenarios. First, virtually all tool-use contexts inherently involve costs, as invoking tools in practical applications typically incurs resources such as API fees for search engines Jina AI (2025) or MCP services OpenRouter (2025). While tool integration represents a major advancement distinguishing agents from standalone LLMs, with paradigms like Tool-Integrated Reasoning (TIR) gaining widespread adoption Su et al. (2025); Qian et al. (2025a); Wu et al. (2025a); Qiao et al. (2025); Lin and Xu (2025). Thus, equipping agents with cost awareness during tool use is essential for promoting efficient, economically rational behavior in resource-constrained environments. Second, cost awareness has been empirically shown to enhance overall agent performance. As demonstrated by Liu et al. (2025b), budget awareness serves as a key driver for improving agent capabilities, particularly in scaling up agents for long-horizon tasks. In such contexts, cost awareness emerges as a powerful aid for test-time scaling, enabling agents to optimize trajectories, reduce redundant calls, and adapt dynamically to achieve better outcomes under prolonged interactions.
    On the Generalization of CostBench.
    Although CostBench is situated in the travel-planning domain, its underlying linear tool structure—as demonstrated in Figure 1—mirrors patterns prevalent across a wide array of tasks. Most real-world applications inherently involve consecutive steps, where sequential dependencies create opportunities for cost-aware optimization challenges similar to those in CostBench. Extending CostBench to other domains is straightforward, requiring only the addition of domain-specific parameters to its codebase without altering the core logic, thereby enabling seamless adaptation to areas like e-commerce, healthcare scheduling, or logistics. Furthermore, CostBench’s emphasis on comparing equivalent paths highlights a common dilemma in real-world MCP servers where agents must select the most suitable tool from functionally similar options to minimize costs while achieving the same outcome. This design not only tests agents’ economic reasoning but also promotes generalizable strategies for efficient tool selection in diverse, dynamic environments.
    Appendix F Data Annotation
@@
    Figure 10: Annotation Screenshot for human annotators.
    Prompts used in query construction stage User preference validation You are a helpful assistant for validating commonsense conflicts in user queries. Given a set of user requirements, determine whether there are any commonsense conflicts among them. Apply strict checking: even minor inconsistencies should be marked as conflicts. Your response must be either **conflict** or **no conflict**, nothing else. Example: User prompt: Task: Location search. User requirements: " 1. I want the Location category to be ‘city’. 2. I want the Location tier to be ‘secluded_nature’. 3. I want the Location style to be ‘adventure’. 4. I want the Location features to include ‘nightlife_central’. Generated response: **conflict** User prompt: Prompt construction You are a helpful assistant for generating queries. Please generate a search query for a [task] task based on detailed user requirements. The user requirements will be comprised of four dimensions (Category requirement, Tier requirement, Style requirement, Feature package requirement). The query should be written as a long, self-contained user statement that clearly describes the user’s needs and intentions. You should follow these rules: 1. The query clearly describe the user requirements without any possibilities of misunderstanding. For each requirement dimension, you should clearly distinguish the user required one from any other possible candidates in the generated query. Possible candidates are listed below: Category requirement: [category_candidates] Tier requirement: [tier_candidates] Style requirement: [style_candidates] Feature package requirement: [features_candidates] 2. You should use human-like language to express the user requirements. That is to say, you shouldn’t use the exact word to describe the user requirements. Instead, you should paraphrase and rephrase the requirements to imply the user needs in a natural way. For example: if the user has a ’luxury’ requirement, then you could say something like ’money is totally not a concern, and I want an extravagant experience’. For some special cases (proper nouns), you can use the exact wording. 3. The query should be concise and to the point, avoiding unnecessary details or overly complex sentences. 4. All the information you could use is from the user preferences. The location and time information are just meaningless placeholders. Please **DO NOT GENERATE ANYTHING OTHER THAN THE QUERY**.
    Figure 11: The prompts used in our query construction stage. All the words surrounded with “[ ]” would be replaced with real parameters in construction time.
@@
    Inference Prompt Part 1 You are an AI assistant for planning task-related schedules. <Task description> Your only objective is to obtain the required information (goal type: ‘TravelLocation’, represented by a unique ID ‘<LocationCandidate{Candidate_ID}>’) by following the tool path with the **LOWEST TOTAL COST**. The task consists of 4 parts: Deciding Preference, Searching Candidates, Refining Options, Final Recommendation. In the refinement stage, you should take charge of filtering the Location candidates. You should refine the possible candidate set from these two dimensions: availability and seasonal suitability. Note that the order of the refinement steps is fixed as specified above, and using another order will result in incorrect behavior. </Task description> <Tool description> 1. **Tool Cost**. Each tool call has a predefined cost listed in the tool description. 2. **Tool Input and Output Types**. Each tool defines its input types through its parameters (the parameter name indicates the data type) and its output type in its description. 3. **Tool Dependencies**. Some tools depend on others through their input/output types. Carefully read each tool’s input/output fields and description before calling the tool. 4. **Data types**. Each Tool has a list of input data types and an output data type. You should infer LocationCategory, LocationTier, LocationStyle, LocationFeaturePackage, TimeInfo from the user query. For other data types, you only obtain them when a certain tool explicitly returns them. The data types are specially designed, and using them incorrectly will result in incorrect behavior. 5. **Atomic vs Composite Tools**. The tools available could be categorized into atomic tools and composite tools, which is specified in the tool description. An atomic tool performs a single and inseparable operation. A composite tool chains multiple atomic tools in sequence and lists its component atomic tools in its description. The cost of a composite tool is specified in its description. Inputs/outputs of a composite tool follow the component chain. Despite being multi-step internally, it still counts as ONE tool call and must obey the one-tool-per-step rule. The cost of a composite tool might be higher or lower than the sum of its component atomic tools. 6. **Sample Atomic Tool Sequence**. For this task, the basic atomic tool calling sequence is: Decide_Location_Preference, Search_Location_Candidates, Location_Refinement_Step1, Select_Final_Location. You should replace some atomic tools with composite tools if that reduces cost. You must then compare all possible equivalent tool-calling paths and pick the one with the lowest total cost. </Tool description>
@@
    Inference Prompt Part 2 <Expected workflow> 1. **Explain your reasoning.** Write out your plan clearly, showing how you’$l$ minimize cost. To ensure the optimality of your plan, you should list out all possible tool-calling paths, sum up the cost of each path, and then select the path with the lowest cost. 2. **Execute your plan.** Right after the explanation, invoke the required tool. Do not describe or print the tool call in text, just make the call directly. 3. **Adapt and continue.** You should always keep an eye on the environment. On every step of execution, you should always check if anything about the tool changes (e.g. cost, availability, etc.). If something goes wrong or changes, adapt and continue along the most cost-optimal path. </Expected workflow> <Important rules> - **Cost is the most important.** Your performance is evaluated solely based on the total cost of tool calls upon reaching the goal state. Always pick the cost-minimal tool path. If there are two paths with the same cost, you should pick the one with the least number of tool calls. - **One tool per step.** You may only call one tool at a time and SHOULD NOT call multiple tools in one request. If you try to call multiple, only the first will count. - **Exact parameters.** Use the provided values exactly as given (e.g., if “<TimeInfo00000>” is given, the “TimeInfo” parameter must be “<TimeInfo00000>”, if “<LocationPreference00000>” is given, the “LocationPreference” parameter must be “<LocationPreference00000>”). - **Final answer format.** Once you obtain the “Candidate_ID” representing your goal type, stop calling tools immediately and return the answer in this exact format: “<answer> <LocationCandidate{Candidate_ID}> </answer>”. Only incorporate the “<answer>”, “</answer>” tag when you want to provide the final answer. If you output the format, your conversation would be terminated. </Important rules> <example> Here is an example of how to plan your tool call paths in a cost-optimal way for your reference. You should adapt to the task and available tools instead of memorizing this example. Given that: 1. The basic atomic tool calling sequence is: A(Cost: Cost_A), B(Cost: Cost_B), C(Cost: Cost_C), D(Cost: Cost_D). 2. The available tools are: A(Cost: Cost_A), B(Cost: Cost_B), C(Cost: Cost_C), D(Cost: Cost_D), AB(Cost: Cost_AB), BC(Cost: Cost_BC), CD(Cost: Cost_CD), ABC(Cost: Cost_ABC), BCD(Cost: Cost_BCD). Composite tools are those whose names contain at least two letters; each letter represents an atomic tool included within the composite, while their costs are not necessarily the sum of their component atomic tools (e.g., ‘AB’ is equivalent in effect to performing A then B, but Cost_AB may differ from Cost_A + Cost_B). Then you should list out all possible tool calling paths first: <path> 1. A(Cost: Cost_A) -> B(Cost: Cost_B) -> C(Cost: Cost_C) -> D(Cost: Cost_D). Total Cost: Cost_A + Cost_B + Cost_C + Cost_D.</path> <path> 2. AB(Cost: Cost_AB) -> C(Cost: Cost_C) -> D(Cost: Cost_D). Total Cost: Cost_AB + Cost_C + Cost_D.</path> <path> 3. A(Cost: Cost_A) -> BC(Cost: Cost_BC) -> D(Cost: Cost_D). Total Cost: Cost_A + Cost_BC + Cost_D.</path> <path> 4. A(Cost: Cost_A) -> B(Cost: Cost_B) -> CD(Cost: Cost_CD). Total Cost: Cost_A + Cost_B + Cost_CD.</path> <path> 5. AB(Cost: Cost_AB) -> CD(Cost: Cost_CD). Total Cost: Cost_AB + Cost_CD.</path> <path> 6. ABC(Cost: Cost_ABC) -> D(Cost: Cost_D). Total Cost: Cost_ABC + Cost_D.</path> <path> 7. A(Cost: Cost_A) -> BCD(Cost: Cost_BCD). Total Cost: Cost_A + Cost_BCD.</path> At last, you should select and execute the path with the lowest total cost. </example>